Ok so here is my problem so far: (substitution)
y=2x+7
3x-y=-9
When you put the y in to substitute should it be like this:
3x-(2x+7)=-9
Would you distribute? Or just totally forget the negative and write it like this:
3x+2x+7=-9

Answer by stanbon(75887)

so far: (substitution)
y=2x+7
3x-y=-9
When you put the y in to substitute should it be like this:
3x-(2x+7)=-9
Would you distribute?
------
Yes:: distribute
=======================
3x-2x-7 = -9
-----
x = -2
--------
Solve for "y":
y = 2x+7
y = -4+7
y = 3
====================
Cheers,
Stan H.
====================